OpenStudy (anonymous):

Fan and Medal Factor 3x^2 – 8x + 4. A. (3x – 4)(x + 1) B. (3x – 4)(x – 1) C. (3x – 2)(x – 2) D. The polynomial is prime

OpenStudy (anonymous):

3x^2-2x-6x+4=x(3x-2)-2(3x-2)=(3x-2)(x-2)